OpenStudy (anonymous):

If f(x)=x^2+3x+5 What Is? f(3+h) :D Help please?

OpenStudy (bibby):

plug in (3+h) everywhere you see x

OpenStudy (anonymous):

solve for h?

OpenStudy (anonymous):

distributive property?

OpenStudy (bibby):

you don't really have to there are some simplifications you do like \((3+h)^2=3*3+3*h+3*h+h^2\) (you can just use (a+b)^2=a^2+2ab+b^2) I assume that's what they want you to do

OpenStudy (bibby):

\( f(x)=x^2+3x+5\longrightarrow f(3+h)=(3+h)^2+3(3+h)+5 \)

OpenStudy (bibby):

yes simplifcation here would involve the distributive property

OpenStudy (bibby):

do you need a little push in the right direction or are you working on it

OpenStudy (anonymous):

kind of i think i'm just messing up the simplification part

OpenStudy (bibby):

show me what you have

OpenStudy (anonymous):

do i apply the distributive property to the beginning f(3+h) so f3+fh?

OpenStudy (bibby):

no, f(blah blah blah) is just a recipe saying t hat the stuff in the parentheses is what replaces the x in our equation

OpenStudy (bibby):

\(f(3+h)=(3+h)^2+3(3+h)+5\) the left hand side doesn't have to be touched, it's just a statement of equivalence

OpenStudy (anonymous):

so im at \[9h^2+9+3h+5\]

OpenStudy (anonymous):

9h is not ^2 my bad

OpenStudy (anonymous):

but i don't think I am supposed to solve for h...

OpenStudy (anonymous):

yes

OpenStudy (bibby):

well, to start \((3+h)^2+3(3+h)+5=\\h^2+6h+9+3(3+h)+5=\\h^2+6h+14+9+3h=?\)

OpenStudy (anonymous):

h^2+9h+23

OpenStudy (anonymous):

thank you :(

OpenStudy (bibby):

looks good